I need these problems worked out. The answer key is included, but I need shown how to do them.

oo Verizon LTE 7:05 PM online.e.u.edu Multiple-choice Questions: Choose the best answer to each question l) You and your friend want to go to the top of the Eiffel Tower. Your friend takes the elevator straight up. You decide to walk up the spiral stairway, taking longer to do so, Compare the gravitational potential energy (U) of you and your friend, after you both reach the top. A) It is impossible to tell, since the times are unknown. B) It is impossible to tell, since the distances are unknown C) Your friends U is greater than your U, because she got to the top faster D) Both of you have the same amount of potential energy E) Your U is greater than your friends U, because you traveled a greater distance in getting to the top 2) A dumbbell-shaped object has two masses of equal mass, m, one at each end of the object. The length of the object from one end to the other is 2R. If this object rotates about an axis perpendicularly through its center, what is its moment of inertia? A) mR 3) A snowboarder coasts on a smooth track that rises from one level to another. If the snowboarders initial speed is 4 m/s, the snowboarder just makes it to the upper level and comes to rest. With a slightly greater initial speed of 5 ms, the snowboarder is still moving to the right on the upper level. What is the snowboarders final speed in this case? A) 1 m/s B) 2 m/s C) 3 m/s D) 4 ms E) 5 m/s 4) A batter hits a 0,140-kg baseball that was approaching him at 30.0 m/s and, as a result, the ball leaves the bat at 40.0 ms in the direction of the pitcher. The ball remains in contact with the bat for 2.0 ms. What is the average force exerted by the bat? (Pay attention to units and direction, or, sign, of momentum.) A) 4900 N B) 4.9 N C) 5300 N D) 700 N 5) A grinding wheel is spinning at a rate of 20.0 revolutions per second. When the power to the grinder is turned off, the grinding wheel slows with constant angular acceleration and takes 80.0 s to come to a rest. What was the angular acceleration of the grinding wheel as it came to rest? (Pay attention to units.) A)157 rad/s, in the same direction as angular velocity B) 1.57 rads, opposite to the direction of angular velocity C) 0.25 rads, in the same direction as angular velocity D) 0.25 rad/s, opposite to the direction of angular velocity
media%2Fa6c%2Fa6c005b1-0130-4923-8898-c7
media%2F256%2F256ccf6d-fe2f-43c9-8af6-b6
media%2F52f%2F52f13dc7-df4a-4d0d-a834-35
Answer #1

1.D

Potential energy = mgh

since potential energy depends on the height and height travelled is same for both.

2. B

I = moment of inertia = mR2 + mR2 = 2 mR2

3.

change in kinetic energy at bottom = kinetic energy at the top

(0.5) m (v22 - v12) = (0.5) m v2

(52 - 42) = v2

v = 3 m/s

4. A

vf = final velocity = 40

vi = initial velocity = - 30

m = mass = 0.140

t = time of contact = 0.002

F = m (vf - vi)/t = (0.140) (40 - (-30))/0.002 = 4900